OpenStudy (staceyg):

Bean plants grow rapidly. Tomorrow a plant will be 23 inches tall, the next day it will be 34 inches tall and on the next day it will be 45 inches tall. Write a rule to represent the height of the bean plant as an arithmetic sequence. How tall will the plant be in 15 days? Please explain this to me as well.

OpenStudy (anonymous):

Well, they grow 11 inches every day. How would you format that into an arithmetic sequence?

OpenStudy (staceyg):

I know that, I don't know how to write it as an arithmetic sequence.

OpenStudy (staceyg):

The options are ; A(n) = 34+(n-1)11;194 inches,A(n)=23+(n-1)11;177 inches,A(n)=13n;195 inches

OpenStudy (staceyg):

@anitasonia98

OpenStudy (anonymous):

You start out with 23 inches, and then add on, so your answer would be 23 + (n-1)11 Now, to find out the amount after 15 days, plug in 15 for n, and solve the equation :)

OpenStudy (staceyg):

I got 407?

OpenStudy (staceyg):

@anitasonia98

OpenStudy (anonymous):

I got a totally different answer. When you plug in 15, you get 14 x 11, which is 154. Add 23 to that, and what do you get?

OpenStudy (staceyg):

Oh I got 177. I was multiplying 23 by 14 then by 11before. Thanks for the help. :) @anitasonia98

OpenStudy (anonymous):

No prob!